OpenStudy (anonymous):

Find the lateral area of a cylinder with base area of 49 π in2 and a height that is twice the radius of its base.

OpenStudy (anonymous):

find the radius of the circular base by solving for r: 49pi = pi*r^2 then with that value of r, your height of the cylinder will be h=2*r

OpenStudy (anonymous):

to find the lateral surface area, just multiply h*(circumference of the circular base) = h*2pi*r

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw...:)